OpenStudy (curry):

integral [0,pi] of (d(theta))/(A-Bcos(theta))

OpenStudy (curry):

\[\int\limits_{0}^{\pi} \frac{ 1 }{ A-Bcos(\theta) }\]

hartnn (hartnn):

try \(\theta = \tan (x/2)\)

hartnn (hartnn):

*** \(x = \tan (\theta/2)\)

hartnn (hartnn):

get cos theta in the form of tan theta/2 find dtheta also

hartnn (hartnn):

its not going to be easy later too, but its a good start

OpenStudy (anonymous):

hartnn's right.

OpenStudy (anonymous):

\[\cos 2x=\frac{ 1-\tan ^2x }{ 1+\tan ^2x }\]

OpenStudy (curry):

Can you guy's simplify it more?

hartnn (hartnn):

t = tan theta/2 dt = 2t/(1+t^2) cos theta = (1-t^2)/(1+t^2) plug all these in the integral few algebraic steps, cancel out 1+t^2 then tell us what you get your new integral in only 't'